OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a, b). 3x + 6y = 3 2x - 10y = 44

OpenStudy (jhannybean):

Multiply the first equation by -2 and multiply the second equation by -3 so we can begin by eliminating the x when we add them together.

OpenStudy (jhannybean):

\[\sf -2(3x+6y=3) \implies\]\[\sf -3(2x-10y=44) \implies \] Work this out first.

OpenStudy (anonymous):

u will get the value of x and y in pairs

OpenStudy (anonymous):

Yeah so 3xy

OpenStudy (jhannybean):

I made a typo in the second post. \[\sf -2(3x+6y=3) \implies\]\[\sf 3(2x-10y=44) \implies\] Work this out instead.